When a current of 2.0 A is supplied to a motor which is used to drive a machine, a mass of 2.0 kg is lifted up for 5.0 m in 2.0 s in a uniform velocity. Assuming that the machine is 100 % efficient, what is the potential difference across the terminals of the motor?


1
Expert's answer
2021-05-30T13:30:50-0400

Mass of the body, "m=2\space kg"

Distance lifted, "s=5\space m"

Time, "t=2\space s"

Work done,"W" = "mgs=2\times10\times5=100\space J"

Power = "\dfrac{W}{t}=\dfrac{100}{2}=50\space J/s"

Let potential difference across the motor be "V"

Current supplied, "I=2\space A"

Power supplied by the motor, "P=VI"

Since the machine is 100 % efficient

"\therefore VI=50"

"V=\dfrac{50}{I}=25\space V"

Potential difference across the terminals of motor = 25 V